# Towards Building Multilingual Language Model for Medicine

### v2

@@ A.0.2 Fine-tuning Prompts
Similar to the prompting technique employed in GPT models, we utilize two distinct types of prompt instructions in our fine-tuning setting so that models can understand whether to output long rationale sentences or not, avoiding confusion when training on the mixture of rationale and multiple-choice data. In detail, for multiple-choice answering, we apply the following prompt: You’re a {language} doctor, kindly address the medical queries according to the patient’s account. Answer with the best option directly. In contrast, to obtain an answer accompanied by its corresponding rationale, we use a more detailed prompt: You’re a {language} doctor, kindly address the medical queries according to the patient’s account in {language}. Let’s solve this step-by-step. You should first give the reason in {language} for your choice. Then you should give the right answer index of the question. It is important to note that during fine-tuning, we employed both of the aforementioned instructions. This implies that for each sample, the model underwent optimization twice in each epoch, with different instructions and corresponding outputs. During evaluation, the first instruction was used to determine the model’s accuracy, owing to its significant convenience in metric calculation and the second instruction is utilized to generate rationales and to assess their similarity with the reference material.
@@ A.0.3 Ranking Prompt
We used the following instructions for rating the rationales with GPT4. Please act as an impartial judge and evaluate the quality of the responses provided by six AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s questions better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Begin your evaluation by comparing the six responses. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. Your output is the ordering of these six models from high to low. Output your final verdict from high to low by strictly following this format: Model A, Model B, Model C, Model D, Model E, and Model F.
Appendix B MMedC Details
B.1 Settings for Filtering Content
@@
In this section, we present the parameters employed for data filtering, which are meticulously configured based on the CulturaX dataset. Our parameterization adopts a relatively stringent criterion aimed at procuring high-quality data.
@@
B.2 Data Statistics
We give the details about the composition of MMedC, including the amount of data coming from different sources for each language.
@@ B.2 Data Statistics
Appendix C Detailed Experiment Results
@@
C.1 Detailed Accuracy on Ablation Study
Here, we give the detailed results of the ablation study, including both accuracy and rationale evaluation across six languages, as shown in Table 6 and 7.
@@ C.1 Detailed Accuracy on Ablation Study
Table 7: Rationale evaluation on MMedB with ROUGE-1/ BLEU-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1.
@@
MMedLM $\mathrm{2}_{en}$ 50.00/45.02 50.81/51.19 50.15/53.11 47.12/43.04 33.06/23.94 47.62/41.64 46.46/42.99
MMedLM $\mathrm{2}_{ocr}$ 49.82/44.75 51.47/52.19 52.66/55.21 47.84/43.68 36.98/28.17 48.93/43.42 47.95/44.57
@@
Table 7: Rationale evaluation on MMedB with ROUGE-1/ BLEU-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1.
C.2 Different Metrics on Rationale Evaluation
@@
We conducted an in-depth analysis of the scores assigned to the rationales generated by various models, utilizing a range of evaluation metrics. Table 8 and  9 illustrate the performance of each model as measured by the N-gram BLEU metrics under PEFT and full fine-tuning settings respectively. Meanwhile, Table 10 and 11 present the scores achieved by these models under the N-gram ROUGE metrics. Table 12 and 13 further show the BERT score of different models, which is a score calculated by averaging N-gram BERT scores as introduced in the main body.
@@
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ BLEU-1 41.45 43.09 28.79 38.89 22.25 42.36 36.14 InternLM BLEU-1 41.29 43.47 22.80 30.35 18.24 36.32 32.08 LLaMA 2 BLEU-1 44.72 42.69 45.58 42.93 31.75 44.22 41.98 ChatDoctor BLEU-1 44.65 40.88 39.54 40.12 30.95 42.88 39.84 MedAlpaca BLEU-1 43.59 40.71 37.27 39.82 30.11 42.80 39.05 PMC-LLaMA BLEU-1 44.98 40.09 38.15 38.89 30.08 43.00 39.20 Mistral BLEU-1 48.13 45.61 43.82 44.73 33.62 47.37 43.88 InternLM 2 BLEU-1 46.87 47.64 42.22 41.81 26.78 44.51 41.64 MMedLM BLEU-1 41.63 44.30 38.61 37.54 19.99 40.79 37.14 MMedLM 2 BLEU-1 47.07 47.15 47.90 43.22 27.81 46.17 43.22 BLOOMZ BLEU-2 14.68 15.03 8.46 14.20 5.04 15.92 12.22 InternLM BLEU-2 14.79 15.28 7.29 9.92 4.59 11.59 10.58 LLaMA 2 BLEU-2 16.29 14.01 18.34 16.57 9.12 16.20 15.09 ChatDoctor BLEU-2 17.46 13.39 14.91 15.98 9.25 16.45 14.57 MedAlpaca BLEU-2 16.81 13.35 13.95 15.85 8.81 16.13 14.15 PMC-LLaMA BLEU-2 17.87 13.27 14.07 14.94 8.71 16.38 14.21 Mistral BLEU-2 20.33 16.26 17.11 18.56 10.56 19.32 17.02 InternLM 2 BLEU-2 18.91 18.53 16.29 15.80 7.22 16.53 15.55 MMedLM BLEU-2 15.33 16.06 14.69 13.76 5.96 15.12 13.49 MMedLM 2 BLEU-2 19.11 18.45 20.09 17.90 8.92 18.98 17.24 BLOOMZ BLEU-3 6.48 6.83 3.50 6.78 1.94 7.50 5.50 InternLM BLEU-3 6.64 6.97 3.48 4.42 1.80 4.93 4.71 LLaMA 2 BLEU-3 7.19 5.91 9.09 8.06 3.74 7.32 6.89 ChatDoctor BLEU-3 8.37 5.82 7.36 7.92 4.00 7.87 6.89 MedAlpaca BLEU-3 8.04 5.80 6.69 7.76 3.68 7.63 6.60 PMC-LLaMA BLEU-3 8.75 5.82 6.81 7.08 3.74 7.82 6.67 Mistral BLEU-3 10.35 7.40 8.57 9.61 4.56 9.68 8.36 InternLM 2 BLEU-3 9.12 8.98 8.10 7.69 2.84 7.83 7.42 MMedLM BLEU-3 6.92 7.48 7.59 6.44 2.33 7.11 6.31 MMedLM 2 BLEU-3 9.21 8.99 10.59 9.21 3.87 9.75 8.60 BLOOMZ BLEU-4 3.25 3.38 1.77 3.26 0.87 4.02 2.76 InternLM BLEU-4 3.53 3.41 1.87 2.14 0.91 2.49 2.39 LLaMA 2 BLEU-4 3.72 2.76 4.94 4.06 1.77 3.85 3.52 ChatDoctor BLEU-4 4.52 2.62 3.99 4.10 1.95 4.37 3.59 MedAlpaca BLEU-4 4.50 2.72 3.60 3.96 1.74 4.18 3.45 PMC-LLaMA BLEU-4 4.90 2.73 3.71 3.46 1.76 4.30 3.48 Mistral BLEU-4 5.92 3.75 4.72 5.18 2.16 5.63 4.56 InternLM 2 BLEU-4 5.07 4.74 4.41 4.07 1.25 4.21 3.96 MMedLM BLEU-4 3.70 3.77 4.27 3.16 1.08 3.87 3.31 MMedLM 2 BLEU-4 5.21 4.71 5.96 5.03 1.93 5.61 4.74 BLOOMZ BLEU 9.95 10.29 5.61 9.91 3.19 11.32 8.38 InternLM BLEU 10.34 10.43 5.33 6.76 2.95 7.79 7.27 LLaMA 2 BLEU 11.14 9.18 13.46 11.69 5.89 11.18 10.42 ChatDoctor BLEU 12.50 8.84 10.94 11.30 6.14 11.87 10.26 MedAlpaca BLEU 12.16 8.93 10.15 11.17 5.76 11.56 9.95 PMC-LLaMA BLEU 13.07 8.91 10.31 10.28 5.80 11.71 10.01 Mistral BLEU 15.10 11.34 12.69 13.59 6.94 14.27 12.32 InternLM 2 BLEU 13.55 13.27 12.05 11.43 4.36 11.76 11.07 MMedLM BLEU 10.66 11.21 11.17 9.51 3.58 10.68 9.47 MMedLM 2 BLEU 13.84 13.16 15.21 13.06 5.73 14.21 12.54

### v3

@@ A.0.2 Fine-tuning Prompts
Similar to the prompting technique employed in GPT models, we utilize two distinct types of prompt instructions in our fine-tuning setting so that models can understand whether to output long rationale sentences or not, avoiding confusion when training on the mixture of rationale and multiple-choice data. In detail, for multiple-choice answering, we apply the following prompt: {mdframed}[backgroundcolor=gray!20] You’re a {language} doctor, kindly address the medical queries according to the patient’s account. Answer with the best option directly. In contrast, to obtain an answer accompanied by its corresponding rationale, we use a more detailed prompt: {mdframed}[backgroundcolor=gray!20] You’re a {language} doctor, kindly address the medical queries according to the patient’s account in {language}. Let’s solve this step-by-step. You should first give the reason in {language} for your choice. Then you should give the right answer index of the question. It is important to note that during fine-tuning, we employed both of the aforementioned instructions. This implies that for each sample, the model underwent optimization twice in each epoch, with different instructions and corresponding outputs. During evaluation, the first instruction was used to determine the model’s accuracy, owing to its significant convenience in metric calculation and the second instruction is utilized to generate rationales and to assess their similarity with the reference material.
@@ A.0.3 Ranking Prompt
We used the following instructions for rating the rationales with GPT4. {mdframed}[backgroundcolor=gray!20] Please act as an impartial judge and evaluate the quality of the responses provided by six AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s questions better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Begin your evaluation by comparing the six responses. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. Your output is the ordering of these six models from high to low. Output your final verdict from high to low by strictly following this format: Model A, Model B, Model C, Model D, Model E, and Model F.
A.1 Rationale Generation Prompt
We used the following instructions for generating rationales for each sample with GPT4. {mdframed}[backgroundcolor=gray!20] You’re a {language} doctor. Analyze the reasons behind choosing this particular option in 100 words for the following question in {language}.
###Question: {question} ###Options: {options} ###Answer: {answer_id}
###Rationale: , where ’{language}’ stands for the language of the question, ’{question}’ stands for the question stem, ’{options}’ stands for the question options and ’{answer}’ stands for the ground-truth answer.
Appendix B Baseline Summarizing
In this section, we conclude all the compared models in Supplementary Table 1. We list out the model size, targeting languages and used training data. For data, we further distinguish them as pre-training data or instruction-tuning data. We use "{model}+{data}" format to denote what base model it used and what data they extra used.
Table 1: A summary table of all our compared existing LLMs. Generally, we split them into three parts, i.e., Close-source LLMs, General Open-source LLMs and Biomedical Open-source LLMs. We report the release time, model size and training data and targeting language types in the table. For training data, we distinguish out whether they further split the data into pre-training data and instruction tuning data because different models may have carried out both steps or only one of them. For Biomedical LLMs since all models are based on a general model and add extra biomedical data for further training, we use "{model}+{data}" format to denote what base model it used and what data they extra used. In table, we use “Unknown” to denote the step exists but the details are confidential and “-” to denote the step is not involved.
Training Data
Model Size Release time Language Types Pre-training Data Instruction Tuning Data
Close-source LLMs
GPT-3.5 Unknown 2022.11 Multilingual Unknown Unknown
GPT-4 Unknown 2023.3 Multilingual Unknown Unknown
Flan-PaLM 540B 2022.12 Multilingual 780B 1836 NLP Tasks
MedPaLM 2 Unknown 2023.5 Multilingual Unknown Unknown
Gemini-1.0 pro Unknown 2024.1 Multilingual Unknown Unknown
General Open-source LLMs
BLOOMZ 7B 2023.5 Multilingual 341B 83 Multilingual NLP Tasks
InternLM 7B 2023.7 Multilingual Unknown Unknown
Llama 2 7B 2023.7 English 2T Unknown
Mistral 7B 2023.10 English Unknown Unknown
InternLM 2 7B 2024.2 Multilingual Unknown Unknown
Llama 3 8B 2024.4 English 15T Unknown
Biomedical Open-source LLMs
ChatDoctor 7B 2023.3 English LLaMA 100K Healthcare Conversation Samples
MedAlpaca 7B 2023.4 English LLaMA 99K Medical QA Samples
PMC-LLaMA 13B 2023.9 English LLaMA + 4B Medical Book Tokens 513K Medical QA or Conversation Samples
MEDITRON 7B 2023.11 English Llama 2 + 46.7B Medical Paper tokens 367K Multiple-Choice QA Samples
BioMistral 7B 2024.2 Multilingual Mistral + 3B PMC Paper Tokens 367K Multiple-Choice QA Samples
MMedLM (Ours) 7B - Multilingual InternLM + MMedC -
MMedLM 2 (Ours) 7B - Multilingual InternLM 2 + MMedC -
MMed-Llama 3 (Ours) 8B - Multilingual Llama 3 + MMedC -
    Since all the English benchmarks share an official test set, we directly use the scores reported in their original papers for other models. For MedAlpaca, GPT-4, GPT-3.5 and Llama 3, the scores are based on Open Medical-LLM Leaderboard [34]
  • In PMC-LLaMA, in addition to book tokens, the authors also sampled a small amount of academical papers as well as the natural corpus to do the mixing for pre-training. For simplicity, in table we only list the mian parts, i.e., 4B book tokens.
Appendix C MMedC Details
C.1 Settings for Filtering Content
@@
In this section, we present the parameters employed for data filtering in Supplementary Table 2, which are meticulously configured based on the CulturaX dataset. Our parameterization adopts a relatively stringent criterion aimed at procuring high-quality data.
[b]
@@
C.2 Data Statistics
We give the details about the composition of MMedC in Supplementary Table 3, including the amount of data coming from different sources for each language.
[b]
@@ C.2 Data Statistics
Appendix D MMedBench Details
We give a detailed statistic of the composition of MMedBench in different languages in Supplementary Table 4.
Table 4: The detailed statistic of MMedBench for different languages. All the token number is reported as a averaged number on all considered cases. We report the train and test statistic number separated as “train/test” in the table. “Question Type” represents the percentage of multiple choices or single choices among on train r test cases.
Language Number of Samples Token Number Question Type
Question Option Rationale Multiple Options Single Option
English 10178/ 1273 181/ 187 38/ 38 208/ 199 0%/ 0% 100%/ 100%
Chinese 27400/ 3426 41/ 42 36/ 36 214/ 209 0%/ 0% 100%/ 100%
French 2171/ 622 27/ 31 71/ 67 207/ 195 72.59%/ 48.39% 27.41%/ 51.61%
Spanish 2657/ 2743 34/ 37 68/ 58 173/ 167 0%/0% 100%/ 100%
Russian 1052/ 256 20/ 20 6/ 6 177/ 173 0%/0% 100%/ 100%
Japanese 1590/ 199 170/ 169 53/ 51 231/ 224 14.28%/ 19.60% 85.72%/ 80.40%
Appendix E Detailed Experiment Results
@@
E.1 Detailed Accuracy on Ablation Study
Here, we give the detailed results of the ablation study, including both accuracy and rationale evaluation across six languages, as shown in Supplementary Table 5 and 6.
@@ E.1 Detailed Accuracy on Ablation Study
Table 6: Rationale evaluation on MMedB with BLEU-1/ ROUGE-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1.
@@
MMedLM $\mathrm{2}_{\text{en}}$ 50.00/45.02 50.81/51.19 50.15/53.11 47.12/43.04 33.06/23.94 47.62/41.64 46.46/42.99
MMedLM $\mathrm{2}_{\text{ocr}}$ 49.82/44.75 51.47/52.19 52.66/55.21 47.84/43.68 36.98/28.17 48.93/43.42 47.95/44.57
@@
MMed-Llama 3
Llama 3 48.74/43.66 49.44/49.42 51.96/53.98 47.11/42.49 34.73/25.07 48.59/42.44 46.76/42.84
[…]
MMed-Llama 3 47.61/42.47 49.96/49.36 52.89/55.06 47.92/42.85 36.32/26.67 48.61/43.32 47.22/43.29
E.2 Different Metrics on Rationale Evaluation
@@
We conducted an in-depth analysis of the scores assigned to the rationales generated by various models, utilizing a range of evaluation metrics. Supplementary Table 7 and Supplementary Table 8 illustrate the performance of each model as measured by the N-gram BLEU metrics under PEFT and full fine-tuning settings respectively. Meanwhile, Supplementary Table 9 and Supplementary Table 10 present the scores achieved by these models under the N-gram ROUGE metrics. Supplementary Table 11 and Supplementary Table 12 further show the BERT score of different models, which is a score calculated by averaging N-gram BERT scores as introduced in the main body.
@@
[ht] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ BLEU-1 41.45 43.09 28.79 38.89 22.25 42.36 36.14 InternLM BLEU-1 41.29 43.47 22.80 30.35 18.24 36.32 32.08 LLaMA 2 BLEU-1 44.72 42.69 45.58 42.93 31.75 44.22 41.98 ChatDoctor BLEU-1 44.65 40.88 39.54 40.12 30.95 42.88 39.84 MedAlpaca BLEU-1 43.59 40.71 37.27 39.82 30.11 42.80 39.05 PMC-LLaMA BLEU-1 44.98 40.09 38.15 38.89 30.08 43.00 39.20 Mistral BLEU-1 48.13 45.61 43.82 44.73 33.62 47.37 43.88 Meditron BLEU-1 44.26 39.26 36.31 38.73 28.34 42.02 38.15 InternLM 2 BLEU-1 46.87 47.64 42.22 41.81 26.78 44.51 41.64 BioMistral BLEU-1 45.85 43.12 38.76 41.82 27.73 45.52 40.46 Llama 3 BLEU-1 46.33 47.09 46.24 43.13 30.89 47.14 43.47 MMedLM (Ours) BLEU-1 41.63 44.30 38.61 37.54 19.99 40.79 37.14 MMedLM 2 (Ours) BLEU-1 47.07 47.15 47.90 43.22 27.81 46.17 43.22 MMed-Llama 3 (Ours) BLEU-1 46.56 47.12 48.10 43.62 33.92 47.67 44.50 BLOOMZ BLEU-2 14.68 15.03 8.46 14.20 5.04 15.92 12.22 InternLM BLEU-2 14.79 15.28 7.29 9.92 4.59 11.59 10.58 LLaMA 2 BLEU-2 16.29 14.01 18.34 16.57 9.12 16.20 15.09 ChatDoctor BLEU-2 17.46 13.39 14.91 15.98 9.25 16.45 14.57 MedAlpaca BLEU-2 16.81 13.35 13.95 15.85 8.81 16.13 14.15 PMC-LLaMA BLEU-2 17.87 13.27 14.07 14.94 8.71 16.38 14.21 Mistral BLEU-2 20.33 16.26 17.11 18.56 10.56 19.32 17.02 Meditron BLEU-2 17.39 13.08 12.06 14.47 7.95 14.97 13.32 InternLM 2 BLEU-2 18.91 18.53 16.29 15.80 7.22 16.53 15.55 BioMistral BLEU-2 18.74 15.15 14.06 16.03 6.45 17.55 14.66 Llama 3 BLEU-2 19.20 17.52 19.16 17.41 8.57 19.02 16.81 MMedLM (Ours) BLEU-2 15.33 16.06 14.69 13.76 5.96 15.12 13.49 MMedLM 2 (Ours) BLEU-2 19.11 18.45 20.09 17.90 8.92 18.98 17.24 MMed-Llama 3 (Ours) BLEU-2 18.99 17.59 21.19 18.37 10.76 20.03 17.82 BLOOMZ BLEU-3 6.48 6.83 3.50 6.78 1.94 7.50 5.50 InternLM BLEU-3 6.64 6.97 3.48 4.42 1.80 4.93 4.71 Llama 2 BLEU-3 7.19 5.91 9.09 8.06 3.74 7.32 6.89 ChatDoctor BLEU-3 8.37 5.82 7.36 7.92 4.00 7.87 6.89 MedAlpaca BLEU-3 8.04 5.80 6.69 7.76 3.68 7.63 6.60 PMC-LLaMA BLEU-3 8.75 5.82 6.81 7.08 3.74 7.82 6.67 Mistral BLEU-3 10.35 7.40 8.57 9.61 4.56 9.68 8.36 Meditron BLEU-3 8.34 5.76 5.74 7.02 3.20 6.62 6.11 InternLM 2 BLEU-3 9.12 8.98 8.10 7.69 2.84 7.83 7.42 BioMistral BLEU-3 9.26 6.76 6.58 7.89 2.28 8.38 6.86 Llama 3 BLEU-3 9.45 8.38 9.85 8.81 3.46 9.53 8.25 MMedLM (Ours) BLEU-3 6.92 7.48 7.59 6.44 2.33 7.11 6.31 MMedLM 2 (Ours) BLEU-3 9.21 8.99 10.59 9.21 3.87 9.75 8.60 MMed-Llama 3 (Ours) BLEU-3 9.43 8.64 11.32 9.60 4.55 10.18 8.95 BLOOMZ BLEU-4 3.25 3.38 1.77 3.26 0.87 4.02 2.76 InternLM BLEU-4 3.53 3.41 1.87 2.14 0.91 2.49 2.39 Llama 2 BLEU-4 3.72 2.76 4.94 4.06 1.77 3.85 3.52 ChatDoctor BLEU-4 4.52 2.62 3.99 4.10 1.95 4.37 3.59 MedAlpaca BLEU-4 4.50 2.72 3.60 3.96 1.74 4.18 3.45 PMC-LLaMA BLEU-4 4.90 2.73 3.71 3.46 1.76 4.30 3.48 Mistral BLEU-4 5.92 3.75 4.72 5.18 2.16 5.63 4.56 Meditron BLEU-4 4.59 2.65 2.91 3.55 1.46 3.42 3.09 InternLM 2 BLEU-4 5.07 4.74 4.41 4.07 1.25 4.21 3.96 BioMistral BLEU-4 5.24 3.26 3.56 4.04 0.96 4.63 3.62 Llama 3 BLEU-4 5.44 4.20 5.61 4.62 1.56 5.38 4.47 MMedLM (Ours) BLEU-4 3.70 3.77 4.27 3.16 1.08 3.87 3.31 MMedLM 2 (Ours) BLEU-4 5.21 4.71 5.96 5.03 1.93 5.61 4.74 MMed-Llama 3 (Ours) BLEU-4 5.43 4.56 6.47 5.37 2.18 5.70 4.95 BLOOMZ BLEU 9.95 10.29 5.61 9.91 3.19 11.32 8.38 InternLM BLEU 10.34 10.43 5.33 6.76 2.95 7.79 7.27 Llama 2 BLEU 11.14 9.18 13.46 11.69 5.89 11.18 10.42 ChatDoctor BLEU 12.50 8.84 10.94 11.30 6.14 11.87 10.26 MedAlpaca BLEU 12.16 8.93 10.15 11.17 5.76 11.56 9.95 PMC-LLaMA BLEU 13.07 8.91 10.31 10.28 5.80 11.71 10.01 Mistral BLEU 15.10 11.34 12.69 13.59 6.94 14.27 12.32 Meditron BLEU 12.47 8.77 8.71 10.25 5.07 10.13 9.23 InternLM 2 BLEU 13.55 13.27 12.05 11.43 4.36 11.76 11.07 BioMistral BLEU 13.62 10.18 10.09 11.48 3.78 12.56 10.29 Llama 3 BLEU 13.97 12.35 14.47 12.53 5.33 14.06 12.12 MMedLM (Ours) BLEU 10.66 11.21 11.17 9.51 3.58 10.68 9.47 MMedLM 2 (Ours) BLEU 13.84 13.16 15.21 13.06 5.73 14.21 12.54 MMed-Llama 3 (Ours) BLEU 13.85 12.74 16.13 13.53 6.91 14.72 12.98